OpenStudy (anonymous):

(2^n+4^n )/(3^n+4^n) is this convergent when n is from 1 to infinity ???????

OpenStudy (anonymous):

i thought some one already posted the answer??

OpenStudy (anonymous):

i solved until it looks like \[((1/2)^{n}+1)/((3/4)^{n} +1)\]

OpenStudy (anonymous):

\[\frac{2^n + 4^n}{3^n + 4^n}\] \[\large{\frac{(\frac{2}{4})^n +1}{(\frac{3}{4})^n +1}}\]

OpenStudy (anonymous):

can anyone tell me how after that step the whole thing becomes convergent??

OpenStudy (anonymous):

If you multiply a fraction by itself it becomes smaller so the ^n terms disappear at infinity leaving 1

OpenStudy (anonymous):

oh..thnx!